OpenStudy (anonymous):

Get a medal :3 What is the value of y so that the ordered pair (2, y) is a solution to the equation 2x + y = -2? A. –6 B. –2 C. -1/2 D. 0

OpenStudy (yanasidlinskiy):

Are you trying to find out what y=??

OpenStudy (anonymous):

@YanaSidlinskiy Yes, look for the value of Y that the ordered pair (2, y) is the solution to the equation 2x + y = -2.

OpenStudy (a1234):

Put 2 in x: 2(2) + y = -2 4 + y = -2 y = -6

OpenStudy (anonymous):

@a1234 Ok, thanks :3

OpenStudy (a1234):

Welcome :).